OpenStudy (anonymous):

solve 2(3^x-2) = 54. would I do 6^x-2 = 54 and then what ?

OpenStudy (anonymous):

I dont get it, I have to find the value for x

OpenStudy (anonymous):

\[2(3^x-2)=54\\3^x-2=27\\3^x=29\\x=\log_b29\]

OpenStudy (anonymous):

... unless you typed it incorrectly.\[2\times3^{x-2}=54\\3^{x-2}=27\\x-2=\log_327=3\\x=5\] As an aside, I meant...\[x=\log_329\]... before.

OpenStudy (anonymous):

I get it :) thank you !